# Dark Matter Search Results from 4.2 Tonne-Years of Exposure of the LUX-ZEPLIN (LZ) Experiment J. Aalbers

### v2

@@ Abstract
We report results of a search for nuclear recoils induced by weakly interacting massive particle (WIMP) dark matter using the LUX-ZEPLIN (LZ) two-phase xenon time projection chamber. This analysis uses a total exposure of tonne-years from 280 live days of LZ operation, of which tonne-years and 220 live days are new. A technique to actively tag background electronic recoils from 214Pb decays is featured for the first time. Enhanced electron-ion recombination is observed in two-neutrino double electron capture decays of 124Xe, representing a noteworthy new background. After removal of artificial signal-like events injected into the data set to mitigate analyzer bias, we find no evidence for an excess over expected backgrounds. World-leading constraints are placed on spin-independent (SI) and spin-dependent WIMP-nucleon cross sections for masses 9 GeV/. The strongest SI exclusion set is cm2 at the 90% confidence level at a mass of 36 GeV/, and the best SI median sensitivity achieved is cm2 for a mass of 40 GeV/.
@@
Astrophysical observations provide strong evidence for the existence of dark matter and its dominance of the matter density of the Universe [1, 2, 3, 4]. Well-motivated as a thermal relic, the weakly interacting massive particle (WIMP) is a leading hypothesis for dark matter, arising in a number of beyond-the-standard-model theories [5, 6, 7]. The LUX-ZEPLIN (LZ) collaboration operates the world’s largest two-phase xenon time projection chamber (TPC), with the primary aim of directly detecting WIMPs through their interactions with xenon nuclei. This technology has produced the most stringent constraints on spin-independent (SI) WIMP-nucleon interactions for 5 GeV/ masses, with all three current-generation experiments having reported first results [8, 9, 10]. In this Letter we present the latest WIMP analysis from LZ, combining a new 220 live-day exposure (WS2024) with the 60 live-day exposure from the first result (WS2022) [8] to perform the most sensitive direct search for WIMPs to date.
@@
This analysis incorporates 220.0 live days collected in the new electrode bias configuration between March 27, 2023 and April 1, 2024. Of this WS2024 live time, 40.9 live days are included from when the detector is in the high-mixing state, due to proximity to injected calibrations, with the remaining 179.1 live days taken in the low-mixing state. The xenon gas pressure throughout the run is 1.860 bar(a) with variation at the level of 0.03%. The level of convective mixing is controlled by adjusting the temperature of the liquid returning to the bottom of the TPC from 177.9 K (high-mixing state) to 174.2 K (low-mixing state), with variations at the level of 0.02% for each. Changing from high-mixing to low-mixing state increases the maximum drift time from 1045 s to 1050 s and reduces the overall liquid height by about 0.2 mm. The consequent 1% increase in S2 signal size is corrected for in analysis.
@@
Event waveforms are processed to find and categorize pulses and interactions. Pulse sizes are normalized by the time-dependent size of single VUV photon pulses measured in each channel, and are reported as the number of photons detected (phd) [16, 17]. Events are categorized as SS if they have a single S1 preceding a single S2. The SS identification efficiency is dependent on the S1 and S2 pulse classification efficiencies. Ionization electrons diffuse as they drift, which can lead to fragmentation of the S2 signal into multiple pulses. The splitting probability is 1% on average across the TPC for S2 pulses relevant to this analysis, as estimated with waveform simulations [18]. For a pulse to be classified as S1, coincident photons must be recorded in at least three PMTs; however, undersized PMT responses and differences in photon transit times can lead to missed S1 pulses. The SS S1 identification efficiency is found to be 90% for all S1 pulse sizes in this analysis, rising to 100% for S1s 35 phd (8 keV). This efficiency is assessed by comparing the event classification of a sample of tritium and deuterium-deuterium (DD) neutron calibration events determined by the event reconstruction algorithms to the corresponding classification obtained from a manual, visual categorization campaign.
@@
Tritium and DD neutrons are among several sources used to calibrate the TPC [19]. Tritium is delivered through radiolabeled methane along with 14C, in an activity ratio of approximately 8:1. The tritiated methane, primarily used to calibrate the ER response, was injected into the TPC while in the high-mixing state to ensure its efficient dispersal and complete removal. Mono-energetic 2.45 MeV neutrons from a collimated DD generator, directed roughly 10 cm below the LXe surface, supply the counterpart for calibrating the NR response. AmLi neutron sources, deployed in source tubes at three different azimuthal positions and three different heights around the TPC, provide additional NR calibration events. Injections of 131mXe and 83mKr at roughly monthly and trimonthly cadences, respectively, in addition to background radon progeny, provide mono-energetic peaks throughout WS2024, used to monitor the stability of the detector response and develop corrections for the normalization of S1 and S2 signals.
@@
S1 and S2 signals exhibit spatial and temporal variations due to factors such as changing detector conditions or non-uniform light collection and electron extraction efficiencies. Corrections are applied to produce standardized values, S1 and S2. An map normalizes the S1 signals to the center of the TPC volume, with correction factors averaging 8% in absolute value. S2 signals are normalized to the radial center of the liquid surface in the plane, in addition to being corrected for the depth-dependent probability of electron attachment to electronegative impurities in the LXe; the average values of these corrections are approximately 13% and 7%, respectively. The responses are also normalized in time, using the tritium calibration period as the reference.
@@
Calibrations also inform the tuning of the nest 2.4.0 [20] model for detector response, in which the scintillation photon and ionization electron gains, and , are key parameters. A scan over viable values for and , performed using tritium data in space and constrained by the energy reconstruction of 131mXe and 83mKr peaks, yields phd/photon and phd/electron. A further tuning to more accurately capture the “ER band,” i.e. the shape of the observed distribution of tritium data, enables nest to reproduce the band median and width to better than 0.2% and 5% in , respectively. The DD data define an analogous “NR band,” for which nest is separately tuned to a similar level of precision via small adjustments in the NR yield parameters. The DD neutron interactions are confined to the top of the detector, where the average electric field is 7% higher than the center. Because the charge yield of 222Rn decays used for spatial corrections is more field-sensitive than NRs, corrections in the DD neutron region overestimate field effects on recombination, leading to a systematic shift in the NR band position as assumed for the full TPC. An upwards shift of 2.4% in S2 is applied to the NR model to compensate for this effect, calculated from electrostatic simulations and nest-based models.
@@
To mitigate analyzer bias, artificial, WIMP-like events are injected at random times into the data pipeline, a process referred to as “salting.” Salt events are pre-generated by combining S1 and S2 pulses from sequestered tritium and AmLi calibration data sets. The pulse sizes for a given event are chosen to best match those expected from an NR energy deposition drawn from a parent distribution consisting of the sum of an exponential (WIMP-like) and a flat spectrum. The salt injection rate and parameters of the parent energy distribution were generated randomly within predefined bounds and kept hidden for the entirety of the analysis. Eight salt events were injected, of which one was removed by analysis cuts, consistent with the NR signal efficiency. The other seven events were revealed as salt following the final definition of the data selection criteria and the likelihood models.
@@
Two sources of background NR events are considered: those originating from neutrino interactions and those from radioactive decays in detector components. The neutrino interactions are from CENS of solar (8B and hep) and cosmic-ray-generated atmospheric neutrinos. Nuclear recoils that originate from spontaneous fission and reactions in detector materials can be vetoed by detecting the scattered neutrons and/or associated rays in the Skin and/or OD. Events vetoed by these systems are used as a sideband that is used to derive an in-situ constraint on the number of neutron events in the WIMP search data. As discussed later, the sideband is incorporated directly into the final statistical analysis; however, a sideband-only fit returns an expectation of SS counts, as shown in Table 1. Secondary estimates from simulations based on detector material radioassays and from data using candidate MS neutron events tagged by the veto detectors return expectations of and neutron counts, respectively, consistent with the primary analysis.
@@
Electron capture (EC) and double electron capture (DEC) decays deposit their energy via X ray and Auger cascades in a more localized profile than a particle of the equivalent energy, leading to higher ionization densities and enhanced electron-ion recombination. As a result, EC and DEC events look more NR-like than decays, as reported in Ref. [25]. Preliminary measurements of this effect for ECs in LZ were performed with sidebands of Skin-tagged and multiple scatter decays of 127Xe and 125Xe, isotopes that have been produced in situ via neutron activation. These yield a charge-yield ratio of for 5.2 keV L-shell EC decays relative to decays of the same energy.
@@
Two-neutrino DEC decays of 124Xe are observed in LZ with a half-life of [26], consistent with that in Ref. [27]. The expected number of 124Xe counts in the ROI is estimated assuming an abundance of [28] and branching ratios for the LL (10.00 keV) and LM (5.98 keV) decay modes of 1.4% and 0.8%, respectively. The LM charge suppression is modeled with the same ratio of 0.87 as used for single EC decays. The overlap of two L-shell cascades in LL-shell capture should result in additional recombination. As an independent measurement of the LL-shell charge yield is not available, the Thomas-Imel (TI) box model [29, 30] is invoked to relate the ionization density to the amount of electron-ion recombination at a given applied electric field. The ratio of the charge yield for LL-shell DEC to decays of the same energy, , is treated as a free parameter in the statistical analysis, bounded on one side by the measured for WS2024 (0.87) and on the other by the result from a preliminary study of increasing the ionization density in the TI box model (0.65).
@@
The largest ER background contribution in this exposure comes from -emitting isotopes in the LXe bulk. Xenon sampling measurements indicate a concentration of ppb Ar g/g and an exposure-averaged concentration of ppq Kr g/g. These are used to estimate the contributions of 39Ar and 85Kr, respectively. Exposure in the high-mixing state includes residual tritium and 14C activity following calibrations, and therefore both their beta spectra are incorporated into the background model for this period.
@@
Radon emanates from materials exposed to xenon, dispersing in the active volume. With a measured FV activity of Bq/kg, 214Pb decays from the 222Rn chain, in particular those to the ground state, form the dominant ER background. A method to identify events with high likelihood of being a 214Pb decay — the radon tag — is developed using simulations of the movement of daughter atoms and ions built from observed flow vectors between 222Rn-218Po -decay pairs. A tag volume in which it is probable to find a 214Pb decay is defined for each identified 218Po alpha decay, co-moving with the projected flow streamline for 81 min or approximately 3 the 214Pb half-life. The radon tag is applied in the low-mixing state, where the flow can be modelled, and when data are available for a continuous 81 min period such that tag volumes can be constructed. For WS2024, the radon tag identifies 15% of the tag-active exposure as 214Pb-rich, containing on average % of the 214Pb, estimated directly by measurements of its excited-state decays. The effective 214Pb activity in the untagged remainder of the tag-active exposure is thus Bq/kg. Both the radon-tagged and untagged data sets are used in the final WIMP analysis; the former can be seen alongside all the WIMP search data in the bottom panel of Fig. 3. The separation into 214Pb-rich and 214Pb-poor exposures offers the benefit of being able to better identify and constrain other background contributions, 124Xe in particular.
@@
Figure 3: Black points show the final set of events passing all selection cuts with the panels distinguishing all events in WS2024 (top) from those that are radon tagged (bottom). The radon-tagged data set contains 60% and 15% of the full FV activity of the 214Pb and dispersed backgrounds, respectively; thus the comparison with all data brings to prominence the non-214Pb backgrounds like 124Xe that are not well-described by the ER band. Dark and light, gray and purple shading show the regions containing 68% and 95% of the ER portion of the background model and a 40 GeV/ WIMP, respectively. Green and orange contours show the same quantiles for the best-fit 124Xe distribution and the accidental coincidence events, respectively. Gray lines show contours of constant ER-equivalent (keVee) and NR-equivalent (keVnr) energy.
@@ 40 GeV/ WIMP
1203
42 0
@@
The data are divided into six mutually exclusive samples, which are simultaneously fit using an unbinned, extended maximum likelihood. Five of these apply to the new 220 live-day exposure and describe: (1) events acquired in the high-mixing state; (2) events acquired in the low-mixing state when the radon tag was inactive; events acquired in the low-mixing state that are (3) radon tagged, and (4) radon untagged; and (5) events tagged by the Skin or OD. The nuisance parameters, which float within their uncertainties, are the expected number of events for each background source, listed in Table 1; the radon tagging efficiency; the neutron veto tagging efficiency; and the
@@
124Xe , bounded between 0.65 and 0.87. Parameters characterizing 222Rn and 220Rn daughter backgrounds, tritium and 14C are independent in the high- and low-mixing state exposures. The sixth sample in the likelihood represents WS2022, containing the data and models from the first, 60 live-day WIMP search result [8].
@@
Figure 4: Comparison between data and the best-fit model for WS2024 corresponding to the values shown in Table 1 for a background plus a 40 GeV/ WIMP signal fit in a projection along the ER band median, normalized to the band width. The 40 GeV/ WIMP component is shown with an amplitude equal to its power-constrained 90% C.L. upper limit. The upper and lower panels select the S1 ranges 3–40 phd and 40–80 phd with values of 0.91 and 0.27, respectively. Shaded bands depict the central interval containing 68% of the combined systematic and statistical uncertainties of the model.
@@
The best-fit number of WIMPs at all masses tested (between 9 GeV/ and 100 TeV/) for the combined WS2024+WS2022 analysis is zero. The goodness of fit of the background-only model is assessed in binned 1D projections of the data and model into S1c, log10(S2c), reconstructed energy, distance from the ER band median, as well as the 2D space of . All samples show excellent model-data agreement, passing a Holm-Bonferroni test [31] with a significance level of 0.05. Figure 4 shows the WS2024 data and model in the 1D projection along the ER band median, illustrating the level of agreement through the tails of the distribution with the inclusion of the modified 124Xe response. The fitted values of the radon tagging efficiency and the 124Xe are consistent across all masses, at % and , respectively, in agreement with the pre-fit expectations.
@@
Figure 5 shows the 90% confidence level upper limit on the spin-independent WIMP-nucleon cross section as a function of mass following inference performed with a two-sided, unbinned profile likelihood ratio test statistic [32]. The limit is power constrained at all WIMP masses to below the median expectation following the recommendation of Ref. [*[][;Hereweemployapowerconstraintwhichusestheexpectedsensitivityoftheexperimentassumingnosignalispresent.]baxter2021recommended]. The fluctuation, and therefore the size of the constraint, is largest between 20 and 50 GeV/. This is a consequence of the inclusion of the WS2022 data from Ref. [8], which exhibit an underfluctuation in this region. The limit for the WS2024 data alone can be found in the Supplemental Material, with an approximate fluctuation across the entire mass range, primarily attributed to an underfluctuation of the accidental coincidence background in the region of largest overlap with expected WIMP distributions. The minimum of the limit curve is cm2 at 36 GeV/ while that of the median expected limit is cm2 at 40 GeV/c2. The best median observation potential is cm2 at 40 GeV/c2. Results of searching for evidence of spin-dependent WIMP-nucleon couplings are discussed in the Appendix.

### v3

@@ Abstract
We report results of a search for nuclear recoils induced by weakly interacting massive particle (WIMP) dark matter using the LUX-ZEPLIN (LZ) two-phase xenon time projection chamber. This analysis uses a total exposure of tonne-years from 280 live days of LZ operation, of which tonne-years and 220 live days are new. A technique to actively tag background electronic recoils from 214Pb decays is featured for the first time. Enhanced electron-ion recombination is observed in two-neutrino double electron capture decays of 124Xe, representing a noteworthy new background. After removal of artificial signal-like events injected into the data set to mitigate analyzer bias, we find no evidence for an excess over expected backgrounds. World-leading constraints are placed on spin-independent (SI) and spin-dependent WIMP-nucleon cross sections for masses 9 GeV/. The strongest SI exclusion set is cm2 at the 90% confidence level and the best SI median sensitivity achieved is cm2, both for a mass of 40 GeV/.
@@
Astrophysical observations provide strong evidence for the existence of dark matter and its dominance of the matter density of the Universe [1, 2, 3, 4]. Well-motivated as a thermal relic, the weakly interacting massive particle (WIMP) is a leading hypothesis for dark matter, arising in a number of beyond-the-standard-model theories [5, 6, 7]. The LUX-ZEPLIN (LZ) collaboration operates the world’s largest two-phase xenon time projection chamber (TPC), with the primary aim of directly detecting WIMPs through their interactions with xenon nuclei. This technology has produced the most stringent constraints on spin-independent (SI) WIMP-nucleon interactions for 5 GeV/ masses, with all three current-generation experiments having reported results [8, 9, 10]. In this Letter we present the latest WIMP analysis from LZ, combining a new 220 live-day exposure (WS2024) with the 60 live-day exposure from the first result (WS2022) [8] to perform the most sensitive direct search for WIMPs to date.
@@
This analysis incorporates 220.0 live days collected in the new electrode bias configuration between March 27, 2023 and April 1, 2024. Of this WS2024 live time, 40.9 live days are included from when the detector is in the high-mixing state, due to proximity to injected calibrations, with the remaining 179.1 live days taken in the low-mixing state. The xenon gas pressure throughout the run is 1.860 bar(a) with variation at the level of 0.03%. The level of convective mixing is controlled by adjusting the temperature of the liquid returning to the bottom of the TPC from 177.9 K (high-mixing state) to 174.2 K (low-mixing state), with variations at the level of 0.02% for each. Changing from high-mixing to low-mixing state increases the maximum drift time from 1045 s to 1050 s and reduces the overall liquid height by about 0.2 mm. The consequent 1% increase in S2 signal size is corrected for in analysis. Stable electron lifetimes
averaging 9.3 ms and 13.8 ms were maintained in the
high-mixing and low-mixing states, respectively.
@@
Event waveforms are processed to find and categorize pulses and interactions. Pulse sizes are normalized by the time-dependent size of single VUV photon pulses measured in each channel, and are reported as the number of photons detected (phd) [16, 17]. Events are categorized as SS if they have a single S1 preceding a single S2. The SS identification efficiency is dependent on the S1 and S2 pulse classification efficiencies. Ionization electrons diffuse as they drift, which can lead to fragmentation of the S2 signal into multiple pulses. The splitting probability is 1% on average across the TPC for S2 pulses relevant to this analysis, as estimated with waveform simulations [18]. For a pulse to be classified as S1, coincident photons must be recorded in at least three PMTs; however, undersized PMT responses and differences in photon transit times can lead to missed S1 pulses. The SS S1 identification efficiency is found to be 90% for all S1 pulse sizes in this analysis, rising to 100% for S1s 35 phd (8 keV). This efficiency is assessed by comparing the event classification of a sample of tritium beta decays and deuterium-deuterium (DD) neutron calibration events determined by the event reconstruction algorithms to the corresponding classification obtained from a manual, visual categorization campaign.
@@
Tritium and DD neutrons are among several sources used to calibrate the TPC [19]. Tritium is delivered through radiolabeled methane along with 14C, in an activity ratio of approximately 8:1. The tritiated methane, primarily used to calibrate the ER response, was injected into the TPC while in the high-mixing state to ensure its efficient dispersal and complete removal by the purification system, with a half-life of 53 hours. Mono-energetic 2.45 MeV neutrons from a collimated DD generator, directed roughly 10 cm below the LXe surface, supply the counterpart for calibrating the NR response. AmLi neutron sources, deployed in source tubes at three different azimuthal positions and three different heights around the TPC, provide additional NR calibration events. Injections of 131mXe and 83mKr at roughly monthly and trimonthly cadences, respectively, in addition to background radon progeny, provide mono-energetic peaks throughout WS2024, used to monitor the stability of the detector response and develop corrections for the normalization of S1 and S2 signals.
@@
S1 and S2 signals exhibit spatial and temporal variations due to factors such as changing detector conditions or non-uniform light collection and electron extraction efficiencies. Corrections are applied to produce standardized values, S1 and S2. An map normalizes the S1 signals to the center of the TPC volume, with correction factors averaging 8% in absolute value. S2 signals are normalized to the radial center of the liquid surface in the plane, in addition to being corrected for the measured electron lifetime; the average values of these corrections are approximately 13% and 7%, respectively. The responses are also normalized in time, using the tritium calibration period as the reference.
@@
Calibrations also inform the tuning of the nest 2.4.0 [20] model for detector response, in which the scintillation photon and ionization electron gains, and , are key parameters. A scan over viable values for and , performed using tritium data in space and constrained by the energy reconstruction of 131mXe and 83mKr peaks, yields phd/photon and phd/electron. A further tuning to more accurately capture the “ER band,” i.e. the shape of the observed distribution of tritium data, enables nest to reproduce the band median and width to better than 0.2% and 5% in , respectively. The DD data define an analogous “NR band,” for which nest is separately tuned to a similar level of precision via small adjustments in the NR yield parameters. For both the ER and NR models, the tuned parameters and the resulting light and charge yields are within the uncertainty bands reported in Ref. [21]. The DD neutron interactions are confined to the top of the detector, where the average electric field is 7% higher than the center. Because the charge yield of 222Rn decays used for spatial corrections is more field-sensitive than NRs, corrections in the DD neutron region overestimate field effects on recombination, leading to a systematic shift in the NR band position as assumed for the full TPC. An upwards shift of 2.4% in S2 is applied to the NR model to compensate for this effect, calculated from electrostatic simulations and nest-based models.
@@
To mitigate analyzer bias, artificial, WIMP-like events are injected at random times into the data pipeline, a process referred to as “salting.” Salt events are pre-generated by combining S1 and S2 pulses from sequestered tritium and AmLi calibration data sets. The pulse sizes for a given event are chosen to best match those expected from an NR energy deposition drawn from a parent distribution consisting of the sum of an exponential (WIMP-like) and a flat spectrum. The salt injection rate and parameters of the parent energy distribution were generated randomly within predefined bounds and kept hidden for the entirety of the analysis. Eight salt events were injected, of which one was removed by analysis cuts, consistent with the NR signal efficiency. The other seven events were revealed as salt following the final definition of the data selection criteria and the likelihood models. More details about the salting distribution can be found in the Supplemental Material [24].
@@
Two sources of background NR events are considered: those originating from neutrino interactions and those from radioactive decays in detector components. The neutrino interactions are from CENS of solar (8B and hep) and cosmic-ray-generated atmospheric neutrinos. Nuclear recoils that originate from spontaneous fission and reactions in detector materials can be vetoed by detecting the scattered neutrons and/or associated rays in the Skin and/or OD. Events vetoed by these systems are used as a sideband to derive an in-situ constraint on the number of neutron events in the WIMP search data. As discussed later, the sideband is incorporated directly into the final statistical analysis; a sideband-only fit returns an expectation of SS counts, as shown in Table 1. Secondary estimates from simulations based on detector material radioassays and from data using candidate MS neutron events tagged by the veto detectors return expectations of and neutron counts, respectively, consistent with the primary analysis.
@@
Electron capture (EC) and double electron capture (DEC) decays deposit their energy via X ray and Auger cascades in a more localized profile than a particle of the equivalent energy, leading to higher ionization densities and enhanced electron-ion recombination. As a result, EC and DEC events look more NR-like than decays, as reported in Ref. [26]. Preliminary measurements of this effect for ECs in LZ were performed with sidebands of Skin-tagged and multiple scatter decays of 127Xe and 125Xe, isotopes that have been produced in situ via neutron activation. These yield a charge-yield ratio of for 5.2 keV L-shell EC decays relative to decays of the same energy [27].
@@
Two-neutrino DEC decays of 124Xe are observed in LZ with a half-life of [28], consistent with that in Ref. [29]. The expected number of 124Xe counts in the ROI is estimated assuming an abundance of [30] and branching ratios for the LL (10.00 keV) and LM (5.98 keV) decay modes of 1.4% and 0.8%, respectively. The LM charge suppression is modeled with the same ratio of 0.87 as used for single EC decays. The overlap of two L-shell cascades in LL-shell capture should result in additional recombination. As an independent measurement of the LL-shell charge yield is not available, the Thomas-Imel (TI) box model [31, 21] is invoked to relate the ionization density to the amount of electron-ion recombination at a given applied electric field. The ratio of the charge yield for LL-shell DEC to decays of the same energy, , is treated as a free parameter in the statistical analysis, bounded on one side by the measured for WS2024 (0.87) and on the other side by the result from a preliminary study of increasing the ionization density in the TI box model (0.65).
@@
The largest ER background contribution in this exposure comes from -emitting isotopes in the LXe bulk. Xenon sampling measurements indicate a concentration of ppb Ar g/g and an exposure-averaged concentration of ppq Kr g/g. These are used to estimate the contributions of 39Ar and 85Kr, respectively. Exposure in the high-mixing state includes residual tritium and 14C activity following their injection for ER calibrations, and both their beta spectra are therefore incorporated into the background model for this period.
@@
Radon emanates from materials exposed to xenon, dispersing in the active volume. With a measured FV activity of Bq/kg, 214Pb decays from the 222Rn chain, in particular those to the ground state, form the dominant ER background. We attribute the higher activity in WS2024 with respect to WS2022 (Bg/kg [25]) to the lower drift field, as charged radon daughter progeny move more slowly out of the fiducial volume. A method to identify events with high likelihood of being a 214Pb decay — the radon tag — is developed using simulations of the movement of daughter atoms and ions built from observed flow vectors between 222Rn-218Po -decay pairs. A tag volume in which it is probable to find a 214Pb decay is defined for each identified 218Po alpha decay, co-moving with the projected flow streamline for 81 min or approximately three times the 214Pb half-life. The tag volume includes two flowpaths to account for neutral and ionized daughters. The radon tag can only be applied in the low-mixing state, where the flow can be modeled. Additional periods in the low-mixing state during which radon tagging is not possible, for example around gaps in data acquisition or after flow disturbances induced by power outages, are labeled as radon-tag-inactive.
To summarize, we have four categories of WS2024 exposure: (1) high-mixing, where a radon tag is not possible; (2) low-mixing, but with the radon tag inactive due to interruptions; (3) low-mixing, active radon tag and radon-tagged, with a high probability to contain 214Pb decays; and (4) low-mixing, active radon tag and radon-untagged, with a lower probability to contain 214Pb decays.
For WS2024, 15% of the low-mixing, tag-active exposure is identified as radon-tagged, containing on average % of the total 214Pb as estimated directly by measurements of its excited-state decays.
The effective 214Pb activity in the untagged remainder of the tag-active exposure is thus Bq/kg. The radon-tagged data set can be seen in the bottom panel of Fig. 3.
The separation into radon-tagged and untagged exposures offers the benefit of being able to better identify and constrain other background contributions, 124Xe in particular.
@@
Figure 3: Black points show the final set of events passing all selection cuts with the panels distinguishing all events in WS2024 (top) from those that are radon tagged (bottom). The radon-tagged data set contains 60% and 15% of the full FV activity of the 214Pb and dispersed backgrounds, respectively; thus, the comparison with all data brings to prominence the non-214Pb backgrounds like 124Xe that are not well-described by the ER band. Dark and light gray and purple shading show the regions containing 68% and 95% of the ER portion of the background model and a 40 GeV/ WIMP, respectively. Green and orange contours show the same quantiles for the best-fit 124Xe distribution and the accidental coincidence events, respectively. Gray lines show contours of constant ER-equivalent (keVee) and NR-equivalent (keVnr) energy.
@@ 40 GeV/ WIMP
1202
410
The data are divided into six mutually exclusive samples, which are simultaneously analyzed using an unbinned, extended maximum likelihood. Five of these apply to the new 220 live-day exposure and describe: the (1-4) WS2024 samples described previously
@@
and (5) events tagged by the Skin or OD. The nuisance parameters, which float within their uncertainties, are the expected number of events for each background source, listed in Table 1; the radon tagging efficiency; the neutron veto tagging efficiency; and the
@@
124Xe , bounded between 0.65 and 0.87. Parameters characterizing 222Rn and 220Rn daughter backgrounds, tritium and 14C are independent in the high- and low-mixing state exposures. The sixth sample in the likelihood represents WS2022, containing the data and models from the first, 60 live-day WIMP search result [8]. The full WS2024 data set is shown in the top panel of Fig. 3. All six samples can be seen in Fig. S1 of the Supplemental Material [24], where we also report their exposures.
@@
Figure 4: Comparison between data and the best-fit model for WS2024 corresponding to the values shown in Table 1 for a background plus a 40 GeV/ WIMP signal fit in a projection along the ER band median, normalized to the band width. The 40 GeV/ WIMP component is shown with an amplitude equal to its power-constrained 90% C.L. upper limit. The upper and lower panels select the S1 ranges 3–40 phd and 40–80 phd with values of 0.55 and 0.59, respectively. Shaded bands depict the central interval containing 68% of the combined systematic and statistical uncertainties of the model.
@@
The best-fit number of WIMPs at all masses tested (between 9 GeV/ and 100 TeV/) for the combined WS2024+WS2022 analysis is zero. The goodness of fit of the background-only model is assessed in binned 1D projections of the data and model into , , reconstructed energy, distance from the ER band median, as well as the 2D space of . All samples show excellent model-data agreement, passing a Holm-Bonferroni test [32] with a significance level of 0.05. Figure 4 shows the WS2024 data and model in the 1D projection along the ER band median, illustrating the level of agreement through the tails of the distribution with the inclusion of the modified 124Xe response. The fitted values of the radon tagging efficiency and the 124Xe are consistent across all masses, at % and , respectively, in agreement with the pre-fit expectations.
@@
Figure 5 shows the 90% confidence level upper limit on the spin-independent WIMP-nucleon cross section as a function of mass following inference performed with a two-sided, unbinned profile likelihood ratio test statistic [33]. The limit is power constrained at all WIMP masses to below the median expectation following the recommendation in Ref. [*[][;Hereweemployapowerconstraintwhichusestheexpectedsensitivityoftheexperimentassumingnosignalispresent.]baxter2021recommended]. The fluctuation, and therefore the size of the constraint, is largest between 20 and 50 GeV/. This is a consequence of the inclusion of the WS2022 data from Ref. [8], which exhibit an underfluctuation in this region. The limit for the WS2024 data alone can be found in the Supplemental Material [24], with an approximate fluctuation across the entire mass range, primarily attributed to an underfluctuation of the accidental coincidence background in the region of largest overlap with expected WIMP distributions. The minimum of the limit curve is cm2 at 40 GeV/ while that of the median expected limit is cm2 at 40 GeV/c2. The best median observation potential is cm2 at 40 GeV/c2. Results of searching for evidence of spin-dependent WIMP-nucleon couplings are discussed in the Appendix.